# How glass breaks – Damage explains the difference between surface and fracture energies in amorphous silica

### v3

@@ Abstract
The difference between free surface energy and fracture toughness in amorphous silica is studied via multi-scale simulations. We combine the homogenization of a molecular dynamics fracture model with a phase-field approach to track and quantify the various energy contributions. We clearly separate free surface energy localized as potential energy on the surface and damage diffusion over a 16-23 Å range around the crack path. The plastic contribution is negligible. These findings, which clarify brittle fracture mechanisms in amorphous silicates, align with toughness measurements in silica.
Surface free energy quantifies the excess energy associated with unbalanced interatomic bonds at a material’s surface. In solids, the bulk state is always energetically more favorable, so creating a surface requires energy input. Griffith [1] postulated a similar concept in his theory of crack propagation, stating that a crack can advance if the elastic energy released exceeds a material-specific threshold. Essentially, both free surface energy and fracture surface energy describe the same underlying phenomenon.
@@
However, in experiments, fracture surface energy is almost always significantly higher. It is the case for polymers [2, 3] and crystalline metals [4] where the difference between free surface and fracture energies are found to span several orders of magnitude, because of plastic dissipation around the crack tip, and also for bulk metallic glasses [5, 6]. One of the rare exceptions is silicon crystals [7, 8, 9], where the difference falls within the precision limits of the methods used, and which can be assumed to break by individual bond rupture at an atomically sharp tip as expected from a brittle material [10],
@@
Intriguingly, for silicate glasses, the archetypes of brittle materials, it has long been known that the measured fracture energy exceeds the surface free energy by a factor of ca. 5 [11, 12, 13]. The difference has been tentatively ascribed to plastic dissipation [14, 11] but as it is difficult to experimentally evidence plasticity at this scale in an amorphous material, let alone quantify it (and also because it openly conflicts with the accepted notion of brittleness), the claim has remained controversial [15] to these days: signs of plasticity at silica crack tips have been discounted [16] almost as soon as found [17]. Worse still, standard models for plastic dissipation in cracks [18] are inapplicable to silicates because their plasticity is essentially devoid of hardening [19].
@@
Faced with this conundrum, we focused on silica glass to investigate the rupture of brittle amorphous silicates at the crossover between the atomic and the continuum scales. Previous atomic scale simulations have recognized the existence of an inelastic zone around the crack tip [20, 21]. Here, we show that we can clearly separate the two contributions to the fracture energy : 1) the energy required for surface formation—accounting for bond breaking and atomic relaxation near the surface—, and 2) the energy due to structural rearrangements extending much deeper into the material, over a distance of approximately 20 Å. We also show that the latter process is best described as damage rather than plasticity or nonlinear elasticity and that the ratio between the two contributions conforms to the experiments. This approach offers a fully consistent framework for understanding fracture in amorphous silica, an archetype of brittle materials.
@@
Simulation methods.—To explore the fracture properties of amorphous silica, the atomistic simulations employed the BKS potential [22] to model atomic interactions and Wolf’s truncation method to handle Coulombic forces [23]. The BKS potential was shown to qualitatively but consistently describe the elastic and plastic mechanical properties of these systems [24, 25]. Amorphous glass samples were generated by randomly distributing atoms within a periodic box. Molecular dynamics simulations were then performed to equilibrate, quench, and test the samples. Athermal deformations, pertinent for systems below the glass transition temperature, were applied through iterative energy minimization using the conjugate gradient method. Further simulation details can be found in the Supplementary Materials.
@@
Diffuse damage.—Fracture simulations were carried out on a 3D sample with dimensions $L_{x}\times L_{y}\times L_{z}=400\times 300\times 100$ Å 3, containing 800k atoms, which was found to be large enough to capture initiation and propagation adequately (Fig. 1). The initial defect, a rounded incision with a radius of $r_{c}=10$Å extends along the $z$ direction on a face normal to the $x$ direction. There is periodicity in the $z$ direction. For the boundary conditions on the $x$ and $y$ faces, a K-field displacement was imposed, simulating uniaxial tension along $y$ at infinity. The crack then initiated and propagated along the $x$ direction as the loading parameter ($K_{I}$) was increased. Further details of the fracture simulation are available in the Supplementary Material.
The atomic scale fields of the cracked sample were then coarse-grained (CG), which essentially involves Gaussian convolution to ensure mass and energy conservation during the discrete-continuum transition. The choice of the CG width $w$ will be discussed in detail below. As a result, the following continuum quantities were available in both Lagrangian and Eulerian configurations: displacements $u$, Hencky (logarithmic) strain tensor $\bm{\varepsilon}$, Cauchy stress tensor $\bm{\sigma}$ , potential energy density $\psi_{\rm{pot}}$ (sum of the atomic interactions) and mass density $\rho$.
@@
$\psi_{\epsilon}\leavevmode\nobreak\ =\leavevmode\nobreak\ {\psi_{\epsilon,0}}+\int_{{\bm{\varepsilon}_{0}}}^{\bm{\varepsilon}}{\bm{\sigma}:d\bm{\varepsilon}}\approx{\psi_{\epsilon,0}}+{\textstyle{1\over 2}}\bm{\sigma}:\bm{\varepsilon}J,$ (1)
To quantify this evolution, we resort to the continuum mechanics concept of damage, where the local material stiffness decreases gradually down to a vanishing stiffness which signals fracture [27]. Damage can thus conveniently be described as a scalar field defined by:
${\psi_{\epsilon}}\left({{\bf{u}},d}\right)={\left({1-d}\right)^{2}}\psi_{0}\left({\bm{\varepsilon}\left({\bf{u}}\right)}\right).$ (2)
Thus $d$ evolves between 0 for the pristine and 1 for the fully damaged material. We have used the spectral energy decomposition scheme [28], i.e., the energy is calculated based on the principal values of the strain tensor, thereby preventing crack opening under compression. However, here, the compressive term is negligible and has been omitted from eq. (2). The detailed calculation can be found in the Supplementary Material.
@@
Fig. 1a illustrates the damage distribution from molecular simulations in an $xy$ plane slice of the simulation box after fracture propagation over ca., 150 Å. Fig. 1b shows a cross section along $y$ at $x=125$ Å with a Gaussian fit from which the damage width $l$ is derived. Finally, Fig. 1c presents the damage width $l$ along the crack. The width varies but typically equals 20 Å along the crack faces. The color map indicates maximum damage values along the $y$ axis. Of course, for consistency, the CG width $w$ must be kept small compared to the width of the damage zone $l$. As shown in detail in the Supplementary Material, we found that for w=35w=3–5 Å, the mean value of $l$ remains the same, with only the fluctuation decreasing. At $w=8$ Å, the mean value starts to increase slightly, but this increase remains within the order of 10 %. As $w$ increases beyond 8 Å, the width of the damage zone, $l$, increases linearly with $w$. The value $w=8$ Å will be used subsequently.
@@
The coarse-grained MD simulations demonstrate that damage around the crack is diffused rather than localized, extending into the material beyond the immediate crack tip. The diffusion width $l$ is notably larger than the CG width used to transition from atomic to continuum scales, indicating that the CG width is sufficient to capture local damage and small enough to expose the actual physical spread of damage. Specifically, for fully opened cracks, the damage zone varies in width between approximately 16 and 23 Å, reflecting a non-uniform distribution of damage that suggests its extent may be influenced by structural heterogeneities such as the ring structure, which governs the brittleness and the differences between short-range and medium-range order in glasses [29]. This local variability suggests that the diffusion length of damage is sufficiently small to be influenced by the underlying structural heterogeneity, and can therefore be regarded as a local material property rather than a global parameter.
Finite element scheme.— To show how the MD results can be fed into a continuum scale description, the damage field was then recalculated using a Finite Element Update (FEMU) scheme on 20 slices of the sample to capture the variation along the thickness. During this process, the fracture toughness $g_{c}$ and a constant damage width $l_{c}$ were fitted to minimize the differences between the damage values obtained from coarse-grained molecular scale simulations and those calculated through a phase-field approach [30]. We assumed that the undamaged strain energy density $\psi_{0}$ is known at the atomic scale and minimized the internal energy of the system expressed as:
where $\Gamma$ represents the crack surface. In the phase-field approach [31], it is approximated using a crack density function ($\gamma_{\Gamma}$):
Details are provided in the Supplementary Material.
@@
Interestingly, the FEMU produced a damage width $l_{c}$ ranging from 12 to 20 Å, depending on the loading state. This result correlates well with the observed coarse-grained damage width, even though $l_{c}$ is treated as a global parameter in the FEMU. This correlation suggests that, while damage diffusion is locally variable, the global parameter $l_{c}$ effectively captures the average behavior of the damage width. When solving the homogeneous phase-field solution [28], a comparable $\approx 2$ nm regularization length was found based on experimental measurements.
In the literature [21], crack length is often calculated based on an assumed position of the crack tip, either from local density or using a singular Williams series fit. These methods are not sufficiently precise to describe the moving crack. Furthermore, they cannot capture multiple crack fronts. Here the FEMU approach provides a first method to determine the crack length which overcomes these drawbacks.
Free surface energy.—The free surface energy was determined using a smaller cubic sample with a side length of 100 Å. We slice the sample in half by removing atomic bonding along a plane, thus replacing the periodic boundary condition in the direction normal to the plane with a free surface. The mechanical equilibrium of this dissected sample was then reached through potential energy minimization. The free surface energy, $2\gamma$, was calculated by dividing the energy difference by the area of the newly exposed surface. Notably, $2\gamma$ was found to be independent of the sample size once it surpassed the 50 Å box length.
We determined $2\gamma=2.8\pm 0.2$ J/m2. This value is somewhat larger than room temperature experimental measurements including humidity ($\approx 0.5$ J/m2) [11, 32, 33] but in agreement with calculated temperature variations [34] and correlates well with simulations using more sophisticated potentials [21].
We observed that after cutting, equilibrium is reached through atomic displacements localized within a few interatomic distances from the surface. This finding aligns well with experimental observations of surface relaxation in crystals, where an exponential decay over the first few atomic layers has been reported [35]. Detailed information regarding the nature of the energy changes is provided in the Supplementary Material.

### v4

@@ Abstract
The relationship between free surface energy and fracture energy in amorphous silica is investigated using multiscale simulations. We combine the homogenization of an atomic scale fracture model with a phase-field approach to track and quantify the different energy contributions. Free surface energy, localized as potential energy at the surface, is clearly separated from damage diffusion extending over a 16-23 Å range around the crack path while the plastic contribution is found negligible. Analysis of the local structure reveals that diffuse damage is primarily associated with changes in the ring structure, whereas variations in free surface energy correlate with changes in the coordination number of silicon atoms. These findings provide new insight into brittle fracture mechanisms in amorphous silicates and are consistent with experimental measurements of silica toughness.
keywords:
Free surface energy , Fracture toughness , Phase-field , Molecular dynamics , Internal length scale
journal: Acta Materialia
[…]
1 Introduction
Surface free energy quantifies the excess energy associated with unbalanced interatomic bonds at a material’s surface. In solids, the bulk state is always energetically more favorable, therefore creating a surface requires an energy input. This quantity is an essential parameter in surface science, as it governs how material surfaces directly influence phenomena such as wetting, adhesion, coating, friction, and corrosion.
On the other hand, Griffith [Griffith1921] postulated a closely related concept that forms the basis of fracture mechanics and is essential for estimating fracture resistance [Mueller2015]. He stated that a free surface can be created and a crack can propagate when the elastic energy released exceeds a material-specific critical value. Essentially, both free surface energy and fracture surface energy describe the same underlying physical phenomenon and in an ideal brittle material, the fracture energy is expected to equal twice the surface energy.
@@ 1 Introduction
However, in experiments, fracture surface energy is almost always significantly larger. It is the case for polymers [Tokuda2015, Smith2022] and crystalline metals [Bikerman1978] where the fracture energy exceeds the free surface energy by several orders of magnitude due to plastic dissipation around the crack tip, as well as for bulk metallic glasses [Xu2010, Yiu2020]. One of the rare exceptions is silicon crystals [Gilman1960, Perez2000, Eaglesham1993], where the difference falls within the precision limits of the methods used, and which can be assumed to break by individual bond rupture at an atomically sharp tip as expected from a brittle material [Gumbsch1995],
@@
Intriguingly, for silicate glasses, the archetypes of brittle materials, it has long been known that the measured fracture energy exceeds the surface free energy by a factor of ca. 5 [Wiederhorn69, Rhee1977, Quinn2017]. The difference has been tentatively ascribed to plastic dissipation [marsh1964plastic, Wiederhorn69] but as it is difficult to experimentally evidence plasticity at this scale in an amorphous material, let alone quantify it (and also because it openly conflicts with the accepted notion of brittleness), the claim has remained controversial [lawn1980atomically] to these days: signs of plasticity at silica crack tips have been discounted [Guin2004] almost as soon as found [Celarie2003]. Worse still, standard models for plastic dissipation in cracks [Wei99] are inapplicable to silicates because their plasticity is essentially devoid of hardening [Kermouche16].
Faced with this conundrum, we focused on silica glass to investigate the rupture of brittle amorphous silicates at the crossover between the atomic and the continuum scales. Previous atomic-scale simulations have recognized the existence of an inelastic zone around the crack tip [Rountree2007, Rimsza2018] and have estimated its size; however, they did not provide a mechanical explanation for its origin, nor did they describe its structural signature.
@@
Here, we show that we can clearly separate the two contributions to the fracture energy : 1) the energy required for surface formation—accounting for bond breaking and atomic relaxation near the surface—, and 2) the energy due to structural rearrangements extending much deeper into the material, over a distance of approximately 20 Å. We also show that the latter process is best described as damage rather than plasticity or nonlinear elasticity and that the relative contributions match the experimental results. This approach offers a fully consistent framework for understanding fracture in amorphous silica, an archetype of brittle materials whose fracture mechanisms exhibit unexpected complexity.
While the macroscopic effect of crack-tip regularization is well known [Irwin1958, Bazant1999, Leguillon2018, Molnar2020TAFMEC], its spatial extent is rarely characterized due to the complexity of experimental observations and the fact that existing studies are mostly focused on highly inhomogeneous materials [Cedolin1983, Brooks2013]. As a consequence, this paper is among the first to not only demonstrate the effect of this diffuse damage, but also to identify its spatial extent and provide a structural explanation for its existence. In doing so, it provides one of the first direct demonstrations of the diffuse damage concept employed in recently popular phase-field fracture techniques [Bourdin2000, Ruffini2015].
2 Methods
@@
To explore the fracture properties of amorphous silica, the atomistic simulations employed the BKS potential [Yuan2001] to model atomic interactions and Wolf’s truncation method to handle Coulombic forces [Carre2007]. The BKS potential was shown to qualitatively but consistently describe the elastic and plastic mechanical properties of these systems [Molnar2016AM, Molnar2017MM]. Amorphous glass samples were generated by randomly distributing atoms within a periodic box. Molecular dynamics simulations were then performed to equilibrate, quench, and test the samples. Athermal (quasi-static) deformations, pertinent for systems below the glass transition temperature, were applied through iterative energy minimization using the conjugate gradient method [Molnar2016AM].
Two types of simulations were conducted: (i) fracture propagation analysis; and (ii) free surface energy calculations.
2.1 Sample generation
The amorphous glass samples were prepared through a random sequential placement of atoms within a periodic simulation box. Following this, molecular dynamics simulations using the LAMMPS software [LAMMPS] were conducted to equilibrate, quench, and test the samples. In this study, we examined two sample sizes: a cubic box with a side length of 100 Å, and a rectangular cuboid with dimensions of $400\times 300\times 100$ Å3. After equilibration at 3000 K, the samples were quenched at a rate of 10 K/ps in an NPT ensemble to a final temperature of 10-5 K. The samples were then deformed in an athermal manner. The sample generation methodology and the potential function employed are detailed in Ref. [Molnar2016JNCS].
2.2 Mechanical deformation
@@
Fracture simulations were carried out on a 3D sample with dimensions $L_{x}\times L_{y}\times L_{z}=400\times 300\times 100$ Å3, containing 800k atoms, which was found to be large enough to capture initiation and propagation adequately. The initial defect, a rounded incision extends along the $z$ direction on a face normal to the $x$ direction. There is periodicity in the $z$ direction. For the boundary conditions on the $x$ and $y$ faces, a K-field displacement was imposed, simulating uniaxial tension along $y$ at infinity. The crack then initiated and propagated along the $x$ direction as the loading parameter ($K_{I}$) was increased.
[…]
$\begin{array}[]{l}{{\hat{u}}_{x}}=\frac{{{K_{I}}}}{{8\mu\pi}}\sqrt{2\pi r}\left[{\left({2\kappa-1}\right)\cos\frac{\theta}{2}-\cos\frac{{3\theta}}{2}}\right],\\ {{\hat{u}}_{y}}=\frac{{{K_{I}}}}{{8\mu\pi}}\sqrt{2\pi r}\left[{\left({2\kappa+1}\right)\sin\frac{\theta}{2}-\sin\frac{{3\theta}}{2}}\right],\end{array}$ (1)
where $r$ and $\theta$ are polar coordinates measured from the initial crack tip, $\mu$ is the shear modulus, $\kappa=3-4\nu$, and $\nu$ is Poisson’s ratio. For the bulk sample, the elastic constants were determined as $\mu=31.7$ GPa and $\nu=0.25$. Note, that the largest displacement increment was smaller than 0.01 Å, which is less than 1 % of the characteristic inter-atomic distance ($r_{Si-O}=1.61$ Å). This ensured that the applied deformation remained independent of the step size. Note, that $H_{\rm{fix}}=15$ Å is larger than the interatomic potential cutoff. A schematic illustration of the initial sample and the deformed configuration is depicted in Fig. 1a and b, with a cut in the $z$ direction. During the progressive advance of the crack tip, $K_{I}$ acts solely as a loading parameter and does not retain its intrinsic physical meaning.
[…]
2.3 Continuum damage
Local continuum quantities were computed using a physically based Gaussian convolution technique, commonly known as the coarse-graining method [Hardy1982, Goldhirsch2002]. This method offers the advantage of conserving both energy and mass during the homogenization procedure. With this technique, homogeneous fields for displacements, displacement gradients, strains, rotations, stresses, and potential energy were obtained for both the initial (Lagrangian) and deformed (Eulerian) configurations. As a result, the following continuum quantities were available in both Lagrangian and Eulerian configurations: displacements $u$, Hencky (logarithmic) strain tensor $\bm{\varepsilon}$, Cauchy stress tensor $\bm{\sigma}$ , potential energy density $\psi_{\rm{pot}}$ (sum of the atomic interactions) and mass density $\rho$.
To execute the convolution, the following function was employed:
$\phi\left(r\right)=\frac{1}{{{w^{3}}{\pi^{3/2}}}}{e^{-\frac{{{r^{2}}}}{{{w^{2}}}}}},$ (2)
where $r$ is the distance between the observation point and the atom, and $w$ is the coarse-graining width. This function is normalized such that its integral in 3D equals 1. More details about the technique can be found in Refs. [Molnar2016JNCS, Molnar2017PRE]. An example of the coarse-grained displacement in the $y$ direction is illustrated in Fig. 1c. The choice of the CG width $w$ will be discussed in detail below.
We recall that from the simulations, the coarse-grained stress, strain, and local initial stiffness fields are available. We can calculate damage in the mechanical sense [Kachanov1958] as the difference between the actual coarse-grained elastic energy ($\psi_{el}$) and the undamaged energy components obtained from the strain field:
${\psi_{el}}\left({{\bf{u}},d}\right)={\left({1-d}\right)^{2}}\psi_{0}^{+}\left({\bm{\varepsilon}\left({\bf{u}}\right)}\right)+\psi_{0}^{-}\left({\bm{\varepsilon}\left({\bf{u}}\right)}\right),$ (3)
with $d$ representing the damage, $\psi_{0}^{+}$ and $\psi_{0}^{-}$ representing the tensile and compressive parts of the undamaged strain energy densities. Note that we assume there is no damage in compression. We use the spectral energy decomposition [Molnar2022EFM] here, because we found a comparable failure surface in crack free deformations than in atomic scale simulations [Molnar2016AM]. We note, that more advanced degradation functions exist and can affect the quantitative value of the phase-field regularization length.
Thus $d$ evolves between 0 for the pristine and 1 for the fully damaged material. This variable represents the progressive loss of stiffness and load-bearing capacity of the material.
$\psi_{\epsilon}~=~{\psi_{\epsilon,0}}+\int_{{\bm{\varepsilon}_{0}}}^{\bm{\varepsilon}}{\bm{\sigma}:d\bm{\varepsilon}}\approx{\psi_{\epsilon,0}}+{\textstyle{1\over 2}}\bm{\sigma}:\bm{\varepsilon}J,$ (4)
The local elastic moduli were calculated by deforming the large sample in 6 elementary ways incrementally by $\delta\epsilon=0.01$ % until 0.5 % strain (3 compression, 3 shear). The local stresses were then correlated to local strains. This method, as used in Ref. [Molnar2016JNCS] to correlate local soft spots to sodium distributions, employs the same numerical parameters. Further details are available in the aforementioned paper.
Damage is calculated in the Lagrangian configuration following principles of solid mechanics, as most quantities are better defined in the initial state. Only Cauchy stresses needed to be interpolated back from the deformed state because coarse-graining atomic pairs in the initial configuration that were no longer in contact often resulted in locally negative strain energy, which is physically impossible. This interpolation involved displacing grid points of the Lagrangian configuration by their coarse-grained displacements, followed by interpolating stress values from the Eulerian grid to these displaced points.
To account for free surfaces in the Eulerian configuration that appear on the crack lips, a correction multiplier was defined based on the ratio of locally interpolated densities from the deformed configuration to those from the initial configuration: $\xi={\rho_{L}}/\rho_{E}^{\text{int}}$. This adjustment compensates for the absence of material at the free surfaces in the deformed configuration, ensuring accurate calculations when parts of the coarse-grained volume are empty.
As amorphous materials lack the ordered structure of crystals, initial local stresses can be found in the quenched material. We assume that these stresses store elastic energy, which is quantified and added to the deformation calculated during loading. The initial strain field can be calculated using linear elasticity as: ${\bm{\varepsilon}_{0}}={\bm{C}^{-1}}{\bm{\sigma}_{0}}$, where $\bm{C}$ is the local rigidity tensor and $\bm{\sigma}_{0}$ is the local initial stress tensor. The initial elastic strain energy density was then obtained by ${\psi_{el,0}}=\frac{1}{2}\bm{\sigma}_{0}:\bm{\varepsilon}_{0}$. We note that this quantity was significantly smaller than the energy from the applied deformation.
2.4 Free surface energy
[…]
Furthermore, after energy minimization—performed in the same manner as for the mechanically deformed sample—local quantities were evaluated. Owing to the originally periodic nature of the problem, coarse-graining was carried out in the Lagrangian configuration, and continuity was assumed at both free surfaces to prevent averaging over the empty volume surrounding the sample.
2.5 Finite element scheme
In finite element formulations, damage is typically obtained by minimizing the total energy of the model with respect to the unknown degrees of freedom, namely the local damage variables. In the case of the phase-field approximation, the material properties—such as the critical fracture toughness ($g_{c}$) and the internal length scale—are assumed to be known. Given the crack-driving force, i.e., the undamaged strain energy density, the governing differential equation can be solved to obtain the local damage field.
In this section, in contrast to the standard implementation, we determine the set of material parameters that best fits the damage field obtained using coarse-grained analysis by minimizing the difference between this field and that computed using a finite element phase-field model.
To show how the discrete results can be fed into a continuum scale description, the damage field was recalculated using a Finite Element Update (FEMU) [Kavanagh1971] scheme on 20 slices of the sample to capture the variation along the thickness. During this process, the fracture toughness $g_{c}$ and a constant damage width $l_{c}$ were fitted to minimize the differences between the damage values obtained from coarse-grained molecular scale simulations and those calculated through a phase-field approach [Bourdin2000]. We assumed that the undamaged strain energy density $\psi_{0}$ is known at the atomic scale and minimized the internal energy of the system expressed as:
where $\Gamma$ represents the crack surface. In the phase-field approach [Pham2011], it is approximated using a crack density function ($\gamma_{\Gamma}$):
The objective of the technique is to update the parameters of a constitutive model so that the results of the phase-field simulation, under appropriate boundary conditions, match as closely as possible the results obtained through atomic scale simulations in the sense of a given norm. We used an AT1 description for the phase-field model with a quadratic degradation function.
The FEMU method utilizes the undamaged tensile energy ($\psi_{0}^{+}$) to obtain the local phase-field damage variable. The approach involves iteratively adjusting the material properties, which are considered homogeneous in this case, to minimize the difference between the damage field obtained from the ratio of damaged to undamaged energies ($\bf{d}_{\rm{MD}}$) and the damage field from finite element calculations ($\bf{d}_{\rm{FEM}}$).
$\bm{\Lambda}={\rm{argmin}}{\left[{{{\bf{d}}_{{\rm{MD}}}}-{{\bf{d}}_{{\rm{FEM}}}}\left(\bm{\Lambda}\right)}\right]^{T}}\left[{{{\bf{d}}_{{\rm{MD}}}}-{{\bf{d}}_{{\rm{FEM}}}}\left(\bm{\Lambda}\right)}\right],$ (7)
[…]
${\bf{M}}d\bm{\Lambda}={\bf{b}},$ (8)
with
$\begin{array}[]{l}{\bf{M}}={\left[{\frac{{\partial{{\bf{d}}_{\rm{FEM}}}}}{{\partial\bm{\Lambda}}}}\right]^{T}}\left[{\frac{{\partial{{\bf{d}}_{\rm{FEM}}}}}{{\partial\bm{\Lambda}}}}\right],\\ {\bf{b}}={\left[{\frac{{\partial{{\bf{d}}_{\rm{FEM}}}}}{{\partial\bm{\Lambda}}}}\right]^{T}}\left[{{{\bf{d}}_{\rm{MD}}}-{{\bf{d}}_{\rm{FEM}}}\left(\bm{\Lambda}\right)}\right].\end{array}$ (9)
The fracture properties were changed until the maximum change in error was smaller than $10^{-6}$. The procedure was executed for 20 equally spaced 2D slices in the $z$ direction.
In the literature [Rimsza2018], crack length is often calculated based on an assumed position of the crack tip, either from local density or using a singular Williams series fit. These methods are not sufficiently precise to describe the moving crack. Furthermore, they cannot capture multiple crack fronts. Here the FEMU approach provides a first method to determine the crack length which overcomes these drawbacks.
As a result, we were able to obtain the crack length, the fracture toughness, and the internal length scale of the material from a single snapshot of deformation without accounting for the boundary conditions, relying solely on local quantities.
3 Results
3.1 Diffuse damage
@@
Fig. 2a illustrates the damage distribution from molecular simulations in an $xy$ plane slice of the simulation box after fracture propagation over ca., 150 Å. Fig. 2b shows a cross section along $y$ at $x=125$ Å with a Gaussian fit from which the damage width $l$ is derived. Finally, Fig. 2c presents the damage width $l$ along the crack. The width varies but typically equals 20 Å along the crack faces. The color map indicates maximum damage values along the $y$ axis. Of course, for consistency, the CG width $w$ must be kept small compared to the width of the damage zone $l$.
As shown in detail in A, we found that for w=35w=3–5 Å, the mean value of $l$ remains the same, with only the fluctuation decreasing. At $w=8$ Å, the mean value starts to increase slightly, but this increase remains within the order of 10 %. As $w$ increases beyond 8 Å, the width of the damage zone, $l$, increases linearly with $w$. Subsequently we use the value $w=8$ Å which is approximately 3 times the size of a SiO4 tetrahedron.
@@
The coarse-grained atomic scale simulations demonstrate that damage around the crack is diffused rather than localized, extending into the material beyond the immediate crack tip. The diffusion width $l$ is notably larger than the CG width used to transition from atomic to continuum scales, indicating that the CG width is sufficient to capture local damage and small enough to expose the actual physical spread of damage. As shown Fig. 2c, the damage diffusion length grows as the crack propagates but ultimately converges to a saturated value where the crack is fully open. Specifically, for fully opened cracks, the damage zone varies in width between approximately 16 and 23 Å, reflecting a non-uniform distribution of damage that suggests its extent may be influenced by structural heterogeneities such as the ring structure, which governs the brittleness and the differences between short-range and medium-range order in glasses [yue2015]. This local variability suggests that the diffusion length of damage is sufficiently small to be influenced by the underlying structural heterogeneity, and can therefore be regarded as a local material property rather than a global parameter.
@@
Interestingly, the FEMU produced a damage width $l_{c}$ ranging from 12 to 20 Å, depending on the loading state. This result correlates well with the observed coarse-grained damage width, even though $l_{c}$ is treated as a global parameter in the FEMU. This correlation suggests that, while damage diffusion is locally variable, the global parameter $l_{c}$ effectively captures the average behavior of the damage width. When solving the homogeneous phase-field solution [Molnar2022EFM], a comparable $\approx 2$ nm regularization length was found based on experimental measurements.
3.2 Local plasticity